# fftvis: A Non-Uniform Fast Fourier Transform Based Interferometric Visibility SimulatorA

### v1

@@ 1 Introduction
The detection and characterization of the Cosmic Dawn and Epoch of Reionization (EoR) through the redshifted 21 cm hyperfine transition of neutral hydrogen represents one of the most promising, yet challenging frontiers in modern cosmology. These epochs mark key phases in the evolution of the universe, signaling the formation of the first stars and galaxies, and the emergence of large-scale structure. Observations of the 21 cm line offer a unique probe of these early times, allowing insights into the astrophysics of reionization, the thermal and ionization history of the intergalactic medium (IGM), and the nature of primordial fluctuations. For reviews of the study of the EoR through the 21 cm line, see Furlanetto et al.; Pritchard & Loeb; Liu & Shaw.
@@
Several ongoing and planned experiments aim to detect this faint signal, including the Murchison Widefield Array (MWA; Tingay et al. 2013; Bowman et al. 2013), the Low Frequency Array (LOFAR; van Haarlem et al. 2013; Patil et al. 2017), the Owens Valley Long Wavelength Array (OVRO-LWA; Eastwood et al. 2019), the Hydrogen Epoch of Reionization Array (HERA; DeBoer et al. 2017; Berkhout et al. 2024), and the Square Kilometer Array (SKA; Santos et al. 2015; Koopmans et al. 2015). Prior work includes also includes the Giant Meter Wave Radio Telescope (GMRT; Paciga et al. 2013) and the Precision Array for Probing the Epoch of Reionization (PAPER; Parsons et al. 2010). However, the 21 cm signal is intrinsically faint, five orders of magnitude weaker in brightness temperature than the dominant astrophysical foregrounds such as Galactic synchrotron emission and extragalactic radio sources. This poses a formidable dynamic range challenge for both instrumentation and data analysis (Liu & Shaw, 2020). Although no experiments have reported a detection of the 21 cm power spectrum, many have put increasingly sensitive upper limits, allowing for rough constraints on the timing of heating in the early Universe (HERA Collaboration et al., 2022a, b, 2023; Mertens et al., 2025; Ghara et al., 2025).
@@
A leading strategy for handling this foreground challenge is to exploit the so-called “EoR window”, a region of Fourier space relatively free from foreground contamination under idealized conditions (Datta et al., 2010; Parsons et al., 2012; Morales et al., 2012; Trott et al., 2012; Hazelton et al., 2013; Liu et al., 2014). The existence of this window relies on the assumption that astrophysical foregrounds are spectrally smooth, while the 21 cm signal is expected to vary rapidly with frequency as it traces fluctuations in the ionization and thermal state of the intertergalactic medium across redshift. Accessing this window requires exquisitely precise calibration, accurate systematics modeling, and foreground mitigation techniques to prevent spectral leakage of foreground power into cosmological modes. The required level of accuracy is extreme—even minor inaccuracies in calibration, instrument modeling, or data processing can introduce frequency-dependent systematics that obscure the cosmological signal. This is especially consequential in interferometric measurements, where chromatic instrument responses couple spatial structure into spectral modes, contaminating the 21 cm power spectrum. Foreground leakage induced by spurious spectral structure is one of the main obstacles to a detection of the 21 cm signal. Overcoming this necessitates not only optimized hardware design, but also rigorous control of analysis pipelines, from calibration, systematics mitigation, and foreground modeling through to power spectrum estimation.
@@
As a result, accurate and efficient simulation tools have become an essential component of all modern 21 cm experiments. By enabling controlled, end-to-end analyses, these tools provide insight into how specific choices impact the final power spectrum measurement. Such simulations allow for a comprehensive validation of specific components of an analysis pipeline, including understanding the influence of various instrumental systematics and their mitigation strategies (Kern et al., 2019; Josaitis et al., 2022; Rath et al., 2024), the consequences of calibration errors (Barry et al., 2016; Ewall-Wice et al., 2017; Byrne et al., 2019; Orosz et al., 2019; Joseph et al., 2020), the robustness of power spectrum estimators (Cheng et al., 2018), the development and validation of new calibration frameworks (Sims et al., 2022; Ewall-Wice et al., 2022; Byrne, 2023; Cox et al., 2024), and the characterization of the effects of radio frequency interference and incomplete data sampling (Wilensky et al., 2020; Pagano et al., 2023; Burba et al., 2024; Chen et al., 2025).
@@
Additionally, the growing adoption of Bayesian inference methods in 21 cm cosmology has highlighted an increasing interest in fast and precise visibility simulations (Burba et al., 2023; Kennedy et al., 2023; Murphy et al., 2024; Zhang et al., 2024; Glasscock et al., 2024; Wilensky et al., 2024). These Bayesian techniques typically focus on constraining individual aspects of the analysis, such as instrumental beam patterns, sky emission models, or systematics, utilizing forward modeling to constrain specific subsets of parameters efficiently. Recently, however, Kern 2025 presented a comprehensive, end-to-end forward modeling simulation framework that directly compares the simulated visibilities to the observed data, significantly increasing computational demands. This full-scale approach highlights an even more pressing need for computationally efficient visibility simulators, capable of supporting extensive parameter exploration within these complex, high-dimensional inference frameworks.
@@
Perhaps most importantly, simulations provide a means to rigorously test and validate analysis pipelines from start to finish. By generating mock observations that incorporate the full complexity of real data, including realistic astrophysical and cosmological models, noise, and instrumental effects, experiments are now able to robustly assess signal loss, calibration bias, and foreground leakage by processing these simulations through to the final power spectrum estimation. Such end-to-end tests are essential for building confidence in upper limits reported (Barry et al., 2019; Mertens et al., 2020; Aguirre et al., 2022; Line et al., 2024). Given the scale of modern radio arrays, such validation efforts often require simulating datasets of immense size and complexity, encompassing thousands of frequency channels and integrations, hundreds of antennas, and sky models with millions of components. As a result, advances in simulator efficiency and scalability are not only beneficial but are necessary for the continued progress and reliability of 21 cm cosmological analyses.
@@
Several existing simulators have been developed to meet these needs, including pyuvsim (Lanman et al., 2019), healvis (Lanman & Kern, 2019), MeqTrees (Noordam & Smirnov, 2010), Montblanc (Perkins et al., 2015), FHD (Sullivan et al., 2012), WODEN (Line, 2022), OSKAR (Dulwich et al., 2009), and RIMEz (Martinot, 2022). Each of these tools offers unique capabilities and trade-offs between computational efficiency, accuracy, and flexibility. For the HERA experiment, the matrix-based visibility simulator, matvis (Kittiwisit et al., 2025), has become the simulator of choice for validation and explorative analysis tasks due to its combination of speed and high accuracy. By decomposing the direct evaluation of visibilities into an outer product of individual antenna responses, matvis significantly improves simulation performance, especially for arrays with a large number of antennas. However, this approach becomes increasingly expensive when applied to the full HERA array, where the sheer number of antennas and the required sky model complexity push the technique’s computational limits.
@@
To address these scaling limitations for simulations of densely-packed, many-element arrays, we explore an alternative approach based on the non-uniform Fast Fourier Transform (NUFFT), which offers a route to accelerate visibility simulations by exploiting the Fourier structure of the radio interferometric measurement equation (RIME). In this paper, we introduce fftvis, a radio interferometric visibility simulator that utilizes the Flatiron Institute’s NUFFT library, finufft (Barnett et al., 2019), for efficient evaluation of visibilities under the point source approximation. Building on the foundation laid by matvis, fftvis emphasizes computational scalability and memory efficiency, while simulating visibilities that match those generated by matvis to near-machine precision. Whenever possible fftvis also matches the existing application programming interface (API) of matvis while offering speed-ups of up to multiple orders of magnitude for simulations of compact, many-element arrays, making it an appealing alternative to matvis in most situations where simulation speed is a bottleneck.
@@ 2.1 Evaluating the Measurement Equation
The fundamental quantity in radio interferometric simulations is the visibility, $\mathbf{V}_{ij}$, which describes the measured interferometric response of a baseline formed by two antennas $i$ and $j$ to incident sky radiation. The Radio Interferometer Measurement Equation (RIME; Hamaker et al. 1996; Smirnov 2011) provides a general expression for this response, linking the observed visibilities to the sky brightness and the instrumental response via a sky integral modulated by a geometric fringe term. In its commonly used Jones-matrix formulation, the RIME for a single time-frequency snapshot is given by (Kohn et al., 2019):
@@
where $\hat{\mathbf{s}}$ is a unit vector on the celestial sphere, $\mathbf{b}_{ij}$ is the baseline vector between the two antennas, $c$ is the speed of light, and $\nu$ is the observing frequency. The antenna beam response, represented by matrix $\mathbf{A}_{i}(\nu,\hat{\mathbf{s}})$ represents the complex, polarized voltage response of antenna $i$ in the direction $\hat{\mathbf{s}}$ at frequency, $\nu$. The sky brightness is described by the coherency matrix $\mathbf{C}(\nu,\hat{\mathbf{s}})$, which, under the assumption of uncorrelated electric field components, takes the form:
@@ 2.2 Non-Uniform Fast-Fourier Transform Evaluation of RIME
Figure 1: Schematic representation of the Flatiron Institute Non-Uniform Fast Fourier Transform (finufft) algorithm as utilized by fftvis. The simulation pipeline consists of five key stages: (Panel A) The user-supplied sky model is weighted by the antenna beam pattern, which is interpolated to each source component’s position. (Panel B) The beam-weighted source intensities are projected onto a fine grid through convolution with a compact gridding kernel $\phi$. (Panel C) The gridded sky model undergoes transformation from image-domain to visibility-domain via Fast Fourier Transform (FFT). (Panel D) The visibility-domain grid is corrected for gridding artifacts through deconvolution with the Fourier transform of the gridding kernel, $\hat{\phi}$. (Panel E) Finally, the algorithm interpolates the gridded visibilities to the exact sampling positions in the $uv$-plane determined by the unique baseline vectors derived from the input antenna positions. This process enables efficient computation of visibilities from arbitrary source distributions and antenna positions while maintaining numerical accuracy within user-specified NUFFT tolerance parameters.
Traditional numerical evaluation of the measurement equation involves directly summing complex exponentials for each baseline-source combination, an approach leading to prohibitive computational complexity, especially for many-element arrays or large sky catalogs. To accelerate the simulation of visibilities under the point source approximation, fftvis utilizes non-uniform Fast Fourier transforms (NUFFTs). By reformulating the RIME as a Fourier-like summation over discretized sky positions, fftvis utilizes the high-performance Type 3 NUFFT implementation provided by the finufft library  [^1] The non-uniform Fast Fourier Transform generalizes the FFT to allow for non-uniform sampling in the input or output domains. The Type 1 NUFFT maps non-uniform spatial data to uniform frequency modes, Type 2 maps uniform spatial data to non-uniform frequencies, and Type 3 transforms between arbitrary non-uniform sources and target points. See Barnett et al. (2019) for full definitions. (Barnett et al., 2019), which enables fast and accurate visibility computations, particularly advantageous for dense interferometric arrays.
@@
$V(\nu)\approx\sum_{j=1}^{N_{\rm sources}}c_{j}\,e^{-2\pi i\nu\mathbf{b}\cdot\mathbf{x}_{j}/c},$ (3)
where $\mathbf{x}_{j}$ represents the non-uniform sky positions of $N_{\rm sources}$ point sources, and $c_{j}$ denotes an element of the $2\times 2$ matrix product $\mathbf{A}^{\dagger}\mathbf{C}\mathbf{A}$. This summation is structurally equivalent to a Type 3 NUFFT, also known as the "non-uniform to non-uniform" transform. In its general form, the Type 3 NUFFT aims to efficiently evaluate the outputs $f_{k}$ at $N$ arbitrary target Fourier modes $\mathbf{s}_{k}\in\mathbb{R}^{d}$ given $N_{\rm sources}$ non-uniform source points $\mathbf{x}_{j}\in\mathbb{R}^{d}$ with complex strengths $c_{j}\in\mathbb{C}$ as follows:
$f_{k}:=\sum_{j=1}^{N_{\rm sources}}c_{j}e^{i\mathbf{s}_{k}\cdot\mathbf{x}_{j}},\quad k=1,\ldots,N_{\rm bls}.$ (4)
In the context of visibility simulation, the target modes $\mathbf{s}_{k}=-2\pi\nu\mathbf{b}_{k}/c$ correspond to baseline coordinates scaled by frequency and therefore vary non-uniformly with both baseline vector ($\mathbf{b}$) and frequency ($\nu$). The source positions $\mathbf{x}_{j}$ are also non-uniformly distributed on the sky. A direct evaluation of the sum in Equation 4 requires $\mathcal{O}(N_{\rm sources}N_{\rm bls})$ operations, which becomes computationally expensive for the large numbers of sources ($N_{\rm sources}$) and baselines ($N_{\rm bls}$) typical in wide-field visibility simulations. To evaluate this sum more efficiently, fftvis uses the finufft library, a state-of-the-art CPU-based NUFFT implementation optimized for speed and numerical accuracy.
The finufft algorithm achieves this speedup via a three-stage process:
@@
    Spreading (Gridding) Step: First, the complex source amplitudes, $c_{j}$, located at their non-uniform positions $\{\mathbf{x}_{j}\}$, are effectively "spread" onto a temporary, intermediate uniform grid. This step involves convolving the sparse source data with a compact, smooth kernel function. finufft utilizes a kernel based on the "exponential of semicircle" function,
@@
[…]
@@
    Deconvolution and Interpolation: Finally, the desired visibility values $f_{k}$ corresponding to the specific non-uniform target baseline coordinates $\{\mathbf{s}_{k}\}$ are extracted from the transformed grid resulting from the FFT. This involves two sub-steps: (a) Deconvolution: The distorting effect of the spreading kernel convolution applied in Step 1 is corrected by dividing the FFT output by the analytic Fourier transform of the gridding kernel. (b) Interpolation: The deconvolved spectrum, still defined on the uniform intermediate grid, is then interpolated to the exact target coordinates $\{\mathbf{s}_{k}\}$ corresponding to the requested baseline vectors and frequencies, yielding the final NUFFT output.
A subtle but essential step in the Type 3 NUFFT is selecting the size of the intermediate uniform grid ($n_{i}$) used for the Fast Fourier Transform stage. Unlike Type 1 and Type 2 transforms, where the grid size is directly related to the number of input or output points, the Type 3 grid must be sufficiently fine to accurately represent the transformation between the potentially wide-ranging non-uniform input source positions $\{\mathbf{x}_{j}\}$ and the non-uniform output baseline coordinates $\{\mathbf{s}_{k}\}$. Specifically, to control aliasing and interpolation errors introduced during the gridding and interpolation steps, this intermediate grid must possess adequate resolution across the full range spanned by both the spatial structure of the sources and the Fourier structure determined by the baselines. For each dimension $i=1,\dots,d$, finufft chooses the grid size $n_{i}$ according to the following:

### v2

@@ 1 Introduction
The detection and characterization of the Cosmic Dawn and Epoch of Reionization (EoR) through the redshifted 21 cm hyperfine transition of neutral hydrogen represents one of the most promising, yet challenging frontiers in modern cosmology. These epochs mark key phases in the evolution of the universe, signaling the formation of the first stars and galaxies, and the emergence of large-scale structure. Observations of the 21 cm line offer a unique probe of these early times, allowing insights into the astrophysics of reionization, the thermal and ionization history of the intergalactic medium (IGM), and the nature of primordial fluctuations. For reviews of the study of the EoR through the 21 cm line, see 2006PhR...433..181F; 2012RPPh...75h6901P; 2020PASP..132f2001L.
@@
Several ongoing and planned experiments aim to detect this faint signal, including the Murchison Widefield Array (MWA; 2013JPhCS.440a2033T; 2013PASA...30...31B), the Low Frequency Array (LOFAR; 2013A&A...556A...2V; 2017ApJ...838...65P), the Owens Valley Long Wavelength Array (OVRO-LWA; 2019AJ....158...84E), the Hydrogen Epoch of Reionization Array (HERA; 2017PASP..129d5001D; 2024PASP..136d5002B), and the Square Kilometer Array (SKA; 2015aska.confE..19S; 2015aska.confE...1K). Prior work also includes the Giant Meter Wave Radio Telescope (GMRT; 2013MNRAS.433..639P) and the Precision Array for Probing the Epoch of Reionization (PAPER; 2010AJ....139.1468P). However, the 21 cm signal is intrinsically faint, five orders of magnitude weaker in brightness temperature than the dominant astrophysical foregrounds such as Galactic synchrotron emission and extragalactic radio sources. This poses a formidable dynamic range challenge for both instrumentation and data analysis (2020PASP..132f2001L). Although no experiments have reported a detection of the 21 cm power spectrum, many have put increasingly sensitive upper limits, allowing for rough constraints on the timing of heating in the early Universe (2022ApJ...924...51A; 2022ApJ...925..221A; 2023ApJ...945..124H; 2025arXiv250305576M; 2025arXiv250500373G).
@@
A leading strategy for handling this foreground challenge is to exploit the so-called “EoR window”, a region of Fourier space relatively free from foreground contamination under idealized conditions (2010ApJ...724..526D; 2012ApJ...756..165P; 2012ApJ...752..137M; 2012ApJ...757..101T; 2013ApJ...770..156H; 2014PhRvD..90b3018L). The existence of this window relies on the assumption that astrophysical foregrounds are spectrally smooth, while the 21 cm signal is expected to vary rapidly with frequency as it traces fluctuations in the ionization and thermal state of the intertergalactic medium across redshift. Accessing this window requires exquisitely precise calibration, accurate systematics modeling, and foreground mitigation techniques to prevent spectral leakage of foreground power into cosmological modes. The required level of accuracy is extreme—even minor inaccuracies in calibration, instrument modeling, or data processing can introduce frequency-dependent systematics that obscure the cosmological signal. This is especially consequential in interferometric measurements, where chromatic instrument responses couple spatial structure into spectral modes, contaminating the 21 cm power spectrum. Foreground leakage induced by spurious spectral structure is one of the main obstacles to a detection of the 21 cm signal. Overcoming this necessitates not only optimized hardware design, but also rigorous control of analysis pipelines, from calibration, systematics mitigation, and foreground modeling through to power spectrum estimation.
@@
As a result, accurate and efficient simulation tools have become an essential component of all modern 21 cm experiments. By enabling controlled, end-to-end analyses, these tools provide insight into how specific choices impact the final power spectrum measurement. Such simulations allow for a comprehensive validation of specific components of an analysis pipeline, including understanding the influence of various instrumental systematics and their mitigation strategies (2019ApJ...884..105K; 2022MNRAS.514.1804J; 2024arXiv240608549R), the consequences of calibration errors (2016MNRAS.461.3135B; 2017MNRAS.470.1849E; 2019ApJ...875...70B; 2019MNRAS.487..537O; 2020MNRAS.492.2017J), the robustness of power spectrum estimators (2018ApJ...868...26C), the development and validation of new calibration frameworks (2021MNRAS.503.2457B; 2022MNRAS.517..910S; 2022MNRAS.517..935S; 2022ApJ...938..151E; 2023ApJ...943..117B; 2024MNRAS.532.3375C), and the characterization of the effects of radio frequency interference and incomplete data sampling (2020MNRAS.498..265W; 2023MNRAS.520.5552P; 2024MNRAS.535..793B; 2025ApJ...979..191C).
@@
Additionally, the growing adoption of Bayesian inference methods in 21 cm cosmology has highlighted an increasing interest in fast and precise visibility simulations (2016ApJS..222....3Z; 2016MNRAS.462.3069S; 2017arXiv170103384S; 2019MNRAS.488.2904S; 2023MNRAS.520.4443B; 2023ApJS..266...23K; 2024MNRAS.534.2653M; 2024MNRAS.530.3412Z; 2024RASTI...3..607G; 2024RASTI...3..400W). These Bayesian techniques typically focus on constraining individual aspects of the analysis, such as instrumental beam patterns, sky emission models, or systematics, utilizing forward modeling to constrain specific subsets of parameters efficiently. Recently, however, 2025arXiv250407090K presented a forward modeling simulation framework that directly compares simulated visibilities against observed data, including a parameterized beam, sky, and 21 cm model as part of the optimization routine, significantly increasing computational demands. This approach highlights an even more pressing need for computationally efficient visibility simulators, capable of supporting extensive parameter exploration within these complex, high-dimensional inference frameworks.
@@
Perhaps most importantly, simulations provide a means to rigorously test and validate analysis pipelines from start to finish. By generating mock observations that incorporate the full complexity of real data, including realistic astrophysical and cosmological models, noise, and instrumental effects, experiments are now able to robustly assess signal loss, calibration bias, and foreground leakage by processing these simulations through to the final power spectrum estimation. Such end-to-end tests are essential for building confidence in upper limits reported (2019PASA...36...26B; 2020MNRAS.493.1662M; 2022ApJ...924...85A; 2024PASA...41...67L). Given the scale of modern radio arrays, such validation efforts often require simulating datasets of immense size and complexity, encompassing thousands of frequency channels and integrations, hundreds of antennas, and sky models with millions of components. As a result, advances in simulator efficiency and scalability are not only beneficial but are necessary for the continued progress and reliability of 21 cm cosmological analyses.
@@
Several existing simulators have been developed to meet these needs, including pyuvsim (2019JOSS....4.1234L), healvis (2019ascl.soft07002L), MeqTrees (2010A&A...524A..61N), Montblanc (2015A&C....12...73P), FHD (2012ApJ...759...17S), WODEN (2022JOSS....7.3676L), OSKAR (2009wska.confE..31D), RIMEz (martinot.2022) and hera_sim [^1] https://github.com/HERA-team/hera_sim. Each of these tools offers unique capabilities and trade-offs between computational efficiency, accuracy, and flexibility. For the HERA experiment, the matrix-based visibility simulator, matvis (2025RASTI...4....1K), has become the simulator of choice for validation and explorative analysis tasks due to its combination of speed and high accuracy. By decomposing the direct evaluation of visibilities into an outer product of individual antenna responses, matvis significantly improves simulation performance, especially for arrays with a large number of antennas. However, this approach becomes increasingly expensive when applied to the full HERA array, where the sheer number of antennas and the required sky model complexity push the technique’s computational limits.
@@
To address these scaling limitations for simulations of densely-packed, many-element arrays, we explore an alternative approach based on the non-uniform Fast Fourier Transform (NUFFT), which offers a route to accelerate visibility simulations by exploiting the Fourier structure of the radio interferometric measurement equation (RIME). In this paper, we introduce fftvis, a radio interferometric visibility simulator that utilizes the Flatiron Institute’s NUFFT library, finufft (barnett2019parallelnonuniformfastfourier), for efficient evaluation of visibilities under the point source approximation. Building on the foundation laid by matvis, fftvis emphasizes computational scalability and memory efficiency, while simulating visibilities that match those generated by matvis to near-machine precision. Whenever possible fftvis also matches the existing application programming interface (API) of matvis while offering speed-ups of up to multiple orders of magnitude for simulations of compact, many-element arrays, making it an appealing alternative to matvis in most situations where simulation speed is a bottleneck.
@@ 2.1 Evaluating the Measurement Equation
The fundamental quantity in radio interferometric simulations is the visibility, $\mathbf{V}_{ij}$, which describes the measured interferometric response of a baseline formed by two antennas $i$ and $j$ to incident sky radiation. The Radio Interferometer Measurement Equation (RIME; 1996A&AS..117..137H; 2011A&A...527A.106S) provides a general expression for this response, linking the observed visibilities to the sky brightness and the instrumental response via a sky integral modulated by a geometric fringe term. In its commonly used Jones-matrix formulation, the RIME for a single time-frequency snapshot is given by (2019ApJ...882...58K):
@@
[…]
@@ 2.2 Non-Uniform Fast-Fourier Transform Evaluation of RIME
Figure 1: Schematic representation of the Flatiron Institute Non-Uniform Fast Fourier Transform (finufft) algorithm as utilized by fftvis. The simulation pipeline consists of five key stages: (Panel A) The user-supplied sky model is weighted by the antenna beam pattern, which is interpolated to each source component’s position. (Panel B) The beam-weighted source intensities are projected onto a fine grid through convolution with a compact gridding kernel $\phi$. (Panel C) The gridded sky model undergoes transformation from image-domain to visibility-domain via Fast Fourier Transform (FFT). (Panel D) The visibility-domain grid is corrected for gridding artifacts through deconvolution with the Fourier transform of the gridding kernel, $\hat{\phi}$. (Panel E) Finally, the algorithm interpolates the gridded visibilities to the exact sampling positions in the uvuv-plane determined by the unique baseline vectors derived from the input antenna positions. This process enables efficient computation of visibilities from arbitrary source distributions and antenna positions while maintaining numerical accuracy within user-specified NUFFT tolerance parameters.
Traditional numerical evaluation of the measurement equation involves directly summing complex exponentials for each baseline-source combination, an approach leading to prohibitive computational complexity, especially for many-element arrays or large sky catalogs. To accelerate the simulation of visibilities under the point source approximation, fftvis utilizes non-uniform Fast Fourier transforms (NUFFTs). By reformulating the RIME as a Fourier-like summation over discretized sky positions, fftvis utilizes the high-performance Type 3 NUFFT implementation provided by the finufft library  [^2] The non-uniform Fast Fourier Transform generalizes the FFT to allow for non-uniform sampling in the input or output domains. The Type 1 NUFFT maps non-uniform spatial data to uniform frequency modes, Type 2 maps uniform spatial data to non-uniform frequencies, and Type 3 transforms between arbitrary non-uniform sources and target points. See barnett2019parallelnonuniformfastfourier for full definitions. (barnett2019parallelnonuniformfastfourier), which enables fast and accurate visibility computations, particularly advantageous for dense interferometric arrays.
@@
$V(\nu)\approx\sum_{j=1}^{N_{\rm sources}}c_{j}\,e^{-2\pi i\nu\mathbf{b}\cdot\mathbf{\hat{s}}_{j}/c},$ (3)
where $\mathbf{s}_{j}$ represents the non-uniform sky positions of $N_{\rm sources}$ point sources, and $c_{j}$ denotes an element of the $2\times 2$ matrix product $\mathbf{A}^{\dagger}\mathbf{C}\mathbf{A}$. This summation is structurally equivalent to a Type 3 NUFFT, also known as the "non-uniform to non-uniform" transform. In its general form, the Type 3 NUFFT aims to efficiently evaluate the outputs $f_{k}$ at $N$ arbitrary target Fourier modes $\mathbf{x}_{k}\in\mathbb{R}^{d}$ given $N_{\rm sources}$ non-uniform source points $\mathbf{s}_{j}\in\mathbb{R}^{d}$ with complex strengths $c_{j}\in\mathbb{C}$ as follows:
$f_{k}:=\sum_{j=1}^{N_{\rm sources}}c_{j}e^{i\mathbf{x}_{k}\cdot\mathbf{s}_{j}},\quad k=1,\ldots,N_{\rm bls}.$ (4)
In the context of visibility simulation, the target modes $\mathbf{x}_{k}=-2\pi\nu\mathbf{b}_{k}/c$ correspond to baseline coordinates scaled by frequency and therefore vary non-uniformly with both baseline vector ($\mathbf{b}$) and frequency ($\nu$). The source positions $\mathbf{s}_{j}$ are also non-uniformly distributed on the sky. A direct evaluation of the sum in Equation 4 requires $\mathcal{O}(N_{\rm sources}N_{\rm bls})$ operations, which becomes computationally expensive for the large numbers of sources ($N_{\rm sources}$) and baselines ($N_{\rm bls}$) typical in wide-field visibility simulations.
The finufft algorithm evaluates this sum extremely efficiently via a three-stage process:
@@
    Spreading (Gridding) Step: First, the complex source amplitudes, $c_{j}$, located at their non-uniform positions $\{\mathbf{s}_{j}\}$, are effectively "spread" onto a temporary, intermediate uniform grid. This step involves convolving the sparse source data with a compact, smooth kernel function. finufft utilizes a kernel based on the "exponential of semicircle" function,
@@
    defined on the interval [1,1][-1,1] where $\beta=2.3w$ is a scalar factor proportional to the size of the gridding kernel. This kernel is selected for its favorable properties, including smoothness and rapid decay in the Fourier domain. The spatial width of the convolution kernel balances the accuracy of the subsequent FFT against the computational cost of the gridding operation. This is achieved by defining a half-width parameter $w$ that is related to the requested precision $\epsilon$, typically via $w=\lceil\log_{10}(1/\epsilon)\rceil+1$, where $\lceil\cdot\rceil$ is the ceiling operator. This convolution step maps the non-uniformly sampled source data onto a regular grid for efficient processing via FFT, while the kernel’s properties help minimize errors introduced during this gridding and subsequent interpolation. The size of this intermediate grid ($N_{\rm grid}$) is also chosen based on $\epsilon$ and the source point and target Fourier mode extents.
@@
    Deconvolution and Interpolation: Finally, the desired visibility values $f_{k}$ corresponding to the specific non-uniform target baseline coordinates $\{\mathbf{x}_{k}\}$ are extracted from the transformed grid resulting from the FFT. This involves two sub-steps: (a) Deconvolution: The distorting effect of the spreading kernel convolution applied in Step 1 is corrected by dividing the FFT output by the analytic Fourier transform of the gridding kernel. (b) Interpolation: The deconvolved spectrum, still defined on the uniform intermediate grid, is then interpolated to the exact target coordinates $\{\mathbf{x}_{k}\}$ corresponding to the requested baseline vectors and frequencies, yielding the final NUFFT output.
The three-stage evaluation of the Type 3 NUFFT (i.e., spreading, FFT, and interpolation) is formally analogous to the “gridding and FFT” procedure traditionally used in interferometric imaging (e.g., 2017isra.book.....T). In both approaches, irregularly sampled data are convolved with a compact kernel to map them onto a uniform grid, transformed by a Fast Fourier Transform, and then interpolated back to the desired coordinates. The distinction lies primarily in the choice and interpretation of the convolution kernel. In finufft, the spreading kernel is selected for its numerical properties – compact support, smoothness, and an analytically tractable Fourier transform – that jointly minimize aliasing in the Fourier domain and enable efficient, vectorized evaluation. The “exponential of semicircle” kernel adopted by finufft provides near-optimal accuracy for a given kernel width while maintaining a small computational footprint, which is the primary source of the algorithm’s efficiency (barnett2019parallelnonuniformfastfourier).
Conceptually, one could instead adopt a physically motivated convolution kernel, such as the Fourier transform of the instrument’s primary beam, to incorporate beam-weighting directly in Fourier space. This choice would make the NUFFT’s gridding operation equivalent to beam-weighted imaging, potentially compact in the $uv$-space for wide-field instruments. However, substituting such a kernel would generally sacrifice the analytic and separable form that enables the high precision and performance of the optimized NUFFT kernel. While the Type 3 NUFFT shares its theoretical foundation with traditional imaging, its efficiency arises from the deliberate use of a mathematically optimized, rather than physically derived, gridding kernel.
A subtle but essential step in the Type 3 NUFFT is selecting the size of the intermediate uniform grid ($n_{i}$) used for the Fast Fourier Transform stage. Unlike Type 1 and Type 2 transforms, where the grid size is directly related to the number of input or output points, the Type 3 grid must be sufficiently fine to accurately represent the transformation between the potentially wide-ranging non-uniform input source positions $\{\mathbf{s}_{j}\}$ and the non-uniform output baseline coordinates $\{\mathbf{x}_{k}\}$. Specifically, to control aliasing and interpolation errors introduced during the gridding and interpolation steps, this intermediate grid must possess adequate resolution across the full range spanned by both the spatial structure of the sources and the Fourier structure determined by the baselines. For each dimension $i=1,\dots,d$, finufft chooses the grid size $n_{i}$ according to the following: